# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ Abstract
As major progress in LLM-based long-form text generation enables paradigms such as retrieval-augmented generation (RAG) and inference-time scaling, safely incorporating private information into the generation remains a critical open question. We present InvisibleInk, a highly scalable long-form text generation framework satisfying rigorous differential privacy guarantees with respect to the sensitive references. It interprets sampling from the LLM’s next-token-distribution as the exponential mechanism over the LLM logits with two innovations. First, we reduce the privacy cost by isolating and clipping only the sensitive information in the model logits (relative to the public logits). Second, we improve text quality by sampling from a small superset of the top-$k$ private tokens. Empirical evaluations demonstrate a consistent 8×8$\times$8 × reduction in computation cost over state-of-the-art baselines to generate long-form private text of the same utility across privacy levels. In summary, InvisibleInk is able to generate private long-form text at less than $10$ times the computation cost of non-private generation.
[…]
@@ 1 Introduction
Differential Privacy (DP) [11] can mitigate such risks by providing provable guarantees that the output distribution is statistically indistinguishable whether or not any single piece of sensitive information was included in the input context. This guarantee is not just desirable but often essential in sensitive domains like healthcare and finance, where information leakage carries severe regulatory (e.g., HIPAA, GDPR) and ethical consequences. Therefore, developing effective DP mechanisms for inference-time generation is critical for the responsible application of LLMs in privacy-sensitive applications. We tackle the challenge of achieving high-fidelity, long-form text generation under the rigorous DP guarantees on the reference text(s).
Figure 1: InvisibleInk interprets differentially private text generation as an iterative application of the exponential mechanism over a subset of the LLM’s clipped logits. Our key innovations are: (a) DClip, an improved clipping function to reduce the sensitivity, and hence, the privacy cost; and (b) Top-$k+$ sampling, a truncated decoding algorithm to improve utility by cleverly selecting a subset of logits to sample each token from.
Despite the importance of this problem, existing approaches for DP text generation suffer from severe practical limitations. Consider the prior state-of-the-art method of Amin et al. [12], which interprets token sampling from the LLM as an instance of the canonical exponential mechanism [13] over the next-token logits. It requires a computational overhead of 100$\times$ that of non-private generation or more, to produce non-degenerate text.111 Generally, the drop in utility from DP can be offset by increasing data and computation, e.g., by averaging over larger batches to reduce sensitivity. (DP model training also admits a similar phenomenon [14].) Thus, we could compare methods by the computation cost necessary to attain given privacy and utility levels. This makes them intractable at scale for large models.
@@
We propose InvisibleInk for DP synthetic text generation. It can produce hundreds of tokens at 8×8$\times$8 × lesser computational cost at the same privacy/utility trade-offs compared to prior SoTA methods.
@@
  • InvisibleInk Framework: We propose InvisibleInk, an exponential mechanism-based [13] sampling framework to allow text generation with rigorous DP guarantees under the replace-by-null adjacency. We make two key innovations: (a) DClip to isolate and clip only the sensitive information in the model logits and avoid paying a privacy cost on prior information on language; and (b) Top-$k+$ decoding: a truncated decoding algorithm that approximates sampling from the top-$k$ private tokens with a small superset of size $k^{\prime}\approx k$, ensuring high utility without losing out on private tokens that are unlikely under a public model.
[…]
@@ 2 Related Work
Generating synthetic data with DP has an extensive literature, particularly over categorical features [20, 21, 22, 23, 24]. Many such approaches rely on the exponential mechanism [25, 26, 27]. Unfortunately, it is cryptographically hard to generate DP synthetic data that preserves all 2-way marginals with polynomial sample complexity [28]; see also [29, Thm. 6.12] and [30, Ch. 9]. However, the advent of generative AI pretrained on web-scale data provides strong priors over data distributions, allowing a partial circumvention of these lower bounds. This led to a flurry of recent work to generate complex image/text data with DP [31, 32, 33, 34, 35, 36]. We focus on practical and scalable DP text generation.
DP text can be generated as a post-processing of DP fine-tuned models [37, 38, 39, 40, 41] or via private inference.333 Private inference can be used on pretrained models prompted with private data (our setting) or non-DP fine-tuned models. Both are conceptually identical as they privatize a set of non-private predictions. We focus on the latter as the private fine-tuning can be prohibitively expensive, especially with large models [14]. Private inference can be further divided based on the type of access to the LLM: via a token/text generation API [42, 36, 41], or through white-box access to model internals such as next-token logits [42, 43, 12]; we focus on the white-box setting. Other DP text manipulation tasks include paraphrasing [44, 45] and next-token prediction [46, 47, 48, 18]. (We adapt [18] to text generation as a baseline in our empirical comparisons.)
@@
The concurrent work of [51] extends the approach of Amin et al. [12] by clustering similar references into a batch. This orthogonal approach can also be integrated with InvisibleInk for synthetic text generation, though we note that clustering may not be applicable to some DP text generation tasks, such as RAG. Additionally, [51] replaces mean aggregation of (clipped) logits with median aggregation, which has a smaller local sensitivity and results in a tighter ex-post DP guarantee [52], i.e., the DP guarantee is evaluated after the algorithm’s output is observed. In contrast, InvisibleInk provides a stronger standard (ex-ante) DP guarantee, which can be evaluated before the algorithm is run. Similar to Amin et al. [12], the DP guarantee in [51] is also data-dependent and shares the same limitations (highlighted at the end of §1).
Prior DP text generation approaches diverge significantly from successful decoding algorithms in the NLP literature. Indeed, sampling from truncated next-token distributions [53, 15, 54, 55, 56, 57] is strongly backed by qualitative [15], quantitative [16, 17], and theoretical [58] evidence; see the survey [8] for further details. InvisibleInk aims to generate high-utility text by adapting truncated non-private decoding to the DP setting.
@@ 3 Preliminaries: Differentially Private Text Generation
Language Models & Decoding. An auto-regressive language model $P$ over a vocabulary $V$ defines a distribution $P(x_{t}|{\bm{x}}_{<t})$ over the next token $x_{t}\in V$ in response to a prefix ${\bm{x}}_{<t}:=(x_{1},\ldots,x_{t-1})\in V^{*}$; here, $V^{*}$ denotes the set of all sequences of elements from $V$. These probabilities are typically obtained as the softmax of the next-token logits $\phi(\cdot\,|\,{\bm{x}}_{<t})$ as $P(x_{t}|{\bm{x}}_{<t})\propto\exp\big{(}\phi(x_{t}|{\bm{x}}_{<t})\big{)}$, where $\propto$” hides the normalizing constant that sums the probabilities to one.
@@
$Q_{\tau}(x_{t}|{\bm{x}}_{<t})\propto\exp\left(\tfrac{1}{\tau}\,\phi(x_{t}|{\bm{x}}_{<t})\right),\,\,\text{and}\,\,\,\,Q_{\text{top-}k}(x_{t}|{\bm{x}}_{<t})\propto\begin{cases}P(x_{t}|{\bm{x}}_{<t})\,,&\text{ if }x_{t}\in V_{k},\\ 0\,,&\text{else}\,,\end{cases}$ (1)
@@
where $V_{k}\subset V$ is the set of $k$ tokens with highest logits $\phi(\,\cdot\,|\,{\bm{x}}_{<t})$. See Fig. 2 for an illustration.
@@
The quality of non-DP long-form text generation is primarily determined by the decoding algorithm and its hyperparameter ($\tau$ or $k$[15, 16]. Typically, these hyperparameters are tuned to balance two competing objectives: reducing (or zeroing out) the probability of generating contextually inappropriate or low-probability tokens, while allowing the selection of moderately probable but contextually appropriate tokens to maintain diversity and coherence. In practice, this typically translates to a temperature $\tau\approx 1$ or slightly smaller, and $k\in[10,100]$ (cf. Fig. 2). Further, truncated decoding approaches such as top-$k$ sampling (or nucleus sampling [15], which uses a slightly different truncation rule) tend to produce the most fluent and coherent text. See the review [8] for details.
@@
Figure 2: Left & Center: Illustration of how two common decoding algorithms—temperature rescaling and top-$k$ sampling—reshape the next-token probabilities. Right: Heatmap of MAUVE scores [16, 17] of synthetic text generated for the MIMIC-IV dataset (without using any sensitive references). Notice that the best generations (highest MAUVE) are obtained at $\tau\approx 1$ and $k\approx 100$; InvisibleInk exhibits similar behavior.
@@
In this work, we wish to generate text with a set of privacy-sensitive references ${\bm{R}}=({\bm{r}}_{1},\ldots,{\bm{r}}_{B})\in(V^{*})^{B}$. This arises in synthetic text generation, where we may instruct the model to generate text similar to given references, or in RAG systems where ${\bm{R}}$ contains retrieved information over confidential knowledge bases. The response sampled from the language model may leak sensitive information from the references ${\bm{R}}$.
Differential Privacy (DP). DP provides formal protections against such privacy leakage. At a high level, a (randomized) text generation algorithm $\mathcal{A}$ satisfies DP at the example level, if for all adjacent reference sets ${\bm{R}},{\bm{R}}^{\prime}$ (that differ in a single sensitive example), the probability distributions $\mathcal{A}({\bm{q}},{\bm{R}})$ and $\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})$ over responses ${\bm{x}}\in V^{*}$ are “nearly indistinguishable”.
@@
A fully specified example-level DP guarantee requires precise notions of adjacency and indistinguishability. We focus on the replace-by-null adjacency, where an adjacent ${\bm{R}}^{\prime}$ is obtained by replacing one element of ${\bm{R}}$ by the empty string (or vice versa); we denote it as ${\bm{R}}\simeq{\bm{R}}^{\prime}$. It is functionally similar to adding or removing a reference, but it simplifies the theoretical technicalities; see [14] or §B for a comparison of adjacency notions. We use the indistinguishability notion of zero-concentrated DP (zCDP) [60]: an algorithm $\mathcal{A}$ satisfies $\rho$-zCDP if
$D_{\alpha}\big{(}\mathcal{A}({\bm{q}},{\bm{R}})\|\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})\big{)}\leq\rho\,\alpha\quad\text{for all}\,\,\alpha>1\,\,\text{ andall adjacent }{\bm{R}}\simeq{\bm{R}}^{\prime}\,,$
@@
where $D_{\alpha}$ denotes the $\alpha$-Rényi divergence. Smaller values of $\rho$ mean that $\mathcal{A}({\bm{q}},{\bm{R}})$ and $\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})$ are closer to each other (in terms of the Rényi divergence). Thus, the example that differs between ${\bm{R}},{\bm{R}}^{\prime}$ has a smaller impact on the output, indicating higher privacy protection. The zCDP guarantee can easily be converted into other notions, e.g., $(\varepsilon,\delta)$-DP [61, 62]. Our techniques translate directly to other notions of adjacency and indistinguishability; see §C for details.
Figure 3: Left two: Histograms of private logits ${\bm{\phi}}_{i}$ and differences from public logits ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ for a synthetic data sample generated from the MIMIC dataset, with $5$th and $95$th percentiles shown by the dotted lines. The spread of values for ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ is significantly smaller (around 10×10$\times$10 ×) than that of ${\bm{\phi}}_{i}$. We find that $\mathsf{DClip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})(y)={\bm{\phi}}_{i}(y)$ for over $95\%$ of all $y\in V$ with $C\approx 1$, while the naive clipping of Amin et al. [12] requires $C\approx 8$. This translates into an 8×8$\times$8 × gain in computational efficiency by allowing smaller batch sizes $B$ to maintain $\tau\approx 1$ (which only depends on the ratio $C/B$). Right two: A small clip norm of $C=1$ introduces significant bias in the method of Amin et al. [12], resulting in a near-uniform distribution over the vocabulary. In contrast, our DClip method (§4) preserves probabilities with minimal distortion even at $C=1$.
@@
DP Text Generation via the Exponential Mechanism. Our goal is to design an algorithm $\mathcal{A}({\bm{q}},{\bm{R}})$ to generate text in response to a given query ${\bm{q}}$ and references ${\bm{R}}$ with a desired $\rho$-zCDP guarantee. We assume the whitebox setting where model logits $\phi(\,\cdot\,|\,{\bm{x}})$ can be queried for any ${\bm{x}}\in V^{*}$. This is in contrast to the API-access setting, where one inference call returns an entire text sequence [36].
@@
We describe the prior state-of-the-art approach of Amin et al. [12], which generates the next token $x_{t}\in V$ via the canonical exponential mechanism [13] over the next-token logits of the given LLM. Given a query ${\bm{q}}$ and sensitive references ${\bm{R}}$, the first step is to obtain the next-token logits ${\bm{\phi}}_{i}=\phi(\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$ for each $i\in[B]$ with an LLM inference call. Second, we clip and aggregate the logits as $\bar{\bm{\phi}}=(1/B)\sum_{i=1}^{B}\mathsf{clip}_{C}({\bm{\phi}}_{i})$, where $\mathsf{clip}_{C}({\bm{\phi}})$ clips each coordinate of ${\bm{\phi}}_{i}$ to lie in $[-C,C]$.444 The clipping employed by Amin et al. [12] incorporates a re-centering step, as logits are invariant to additive shifts. While it is shown to be empirically important, we omit this detail in our exposition as it does not change its mathematical properties. We can then sample a $\rho$-zCDP token $x_{t}\in V$ by sampling [63]
@@
where the sensitivity $\mathsf{sens}(\bar{\bm{\phi}})$ measures the maximum change in $\bar{\bm{\phi}}(y)$, when we move to an adjacent set ${\bm{R}}^{\prime}\simeq{\bm{R}}$ of references for any $y\in V$. The clipping and aggregating steps control the sensitivity as $\mathsf{sens}(\bar{\bm{\phi}})=O(C/B)$,555 We have $\mathsf{sens}(\bar{\bm{\phi}})=2C/B$ (resp. $C/B$) under replace-by-null (resp. zero-out) adjacency; cf. §B. We give this method an advantage in §5 by (incorrectly) taking $\mathsf{sens}(\bar{\bm{\phi}})=C/B$ under replace-by-null adjacency. so that Eq. 2 reduces to sampling with temperature $\tau=O(C/B\sqrt{\rho})$.
The clip norm $C$ is a tunable parameter. If it is too small, the resulting distortions in the next-token logits boost the probabilities of the unlikely tail tokens, see Fig. 3. On the other hand, if $C$ is too large, the temperature $\tau\propto C/B$ becomes too large. Both scenarios lead to degenerate text.
Thus, generating non-degenerate text from this method requires taking $C$ sufficiently large (e.g. [12] recommend $C\approx 10$), and $B$ large enough to ensure a temperature of $\tau=O(C/B\sqrt{\rho})\approx 1$. That is, generating each private token requires a large number $B$ of LLM next-token inference calls.666 In practice, we can typically control $B$, as this can be the number of reference documents we use for one synthetic generation, or the number of retrieved references in RAG systems. We find in §5 that this method fails to produce coherent text at $B<O(50)$.

### v2

@@ Abstract
As major progress in LLM-based long-form text generation enables paradigms such as retrieval-augmented generation (RAG) and inference-time scaling, safely incorporating private information into the generation remains a critical open question. We present InvisibleInk, a highly scalable long-form text generation framework satisfying rigorous differential privacy guarantees with respect to the sensitive reference texts. It interprets sampling from the LLM’s next-token-distribution as the exponential mechanism over the LLM logits with two innovations. First, we reduce the privacy cost by isolating and clipping only the sensitive information in the model logits (relative to the public logits). Second, we improve text quality by sampling without any privacy cost from a small superset of the top-$k$ private tokens. Empirical evaluations demonstrate a consistent $8\times$ (or more) reduction in computation cost over state-of-the-art baselines to generate long-form private text of the same utility across privacy levels. InvisibleInk is able to generate, for the first time, high-quality private long-form text at less than $4\text{-}8\times$ times the computation cost of non-private generation, paving the way for its practical use. We open-source a pip-installable Python package (invink) for InvisibleInk at https://github.com/cerai-iitm/invisibleink.
@@ 1 Introduction
Differential Privacy (DP) [11] can mitigate such risks by providing provable guarantees that the output distribution is statistically indistinguishable whether or not any single piece of sensitive information was included in the input context. This guarantee is not just desirable but often essential in sensitive domains like healthcare and finance, where information leakage carries severe regulatory (e.g., HIPAA, GDPR) and ethical consequences and merely anonymizing personally identifiable information (PII) is not sufficient [12] Therefore, developing effective DP mechanisms for inference-time generation is critical for the responsible application of LLMs in privacy-sensitive applications. We tackle the challenge of achieving high-fidelity, long-form text generation under the rigorous DP guarantees on sensitive source text(s) used as references for generation.
Figure 1: InvisibleInk interprets differentially private text generation as an iterative application of the exponential mechanism over a subset of the LLM’s clipped logits. Our key innovations are: (a) DClip, an improved clipping function to reduce the sensitivity, and hence, the privacy cost; and (b) Top-$k+$ sampling, a truncated decoding algorithm to improve utility by selecting a subset of logits to sample each token from.
[…]
@@
We propose InvisibleInk for DP synthetic text generation. It can produce hundreds of tokens at $8\times$ lower computational cost at the same privacy/utility trade-offs compared to prior SoTA methods.
@@
  • InvisibleInk Framework: We propose InvisibleInk, an exponential mechanism-based [14] sampling framework to allow LLM-based text generation with rigorous DP guarantees under the replace-by-null adjacency. We make two key innovations: (a) DClip to isolate and clip only the sensitive information in the model logits and avoid paying a privacy cost on prior information on language; and (b) Top-$k+$ decoding: a truncated decoding algorithm that approximates sampling from the top-$k$ private tokens with a small superset of size $k^{\prime}\approx k$, ensuring high utility without losing out on private tokens that are unlikely under a public model.
  • Privacy-Utility-Compute Tradeoffs: By empirically analyzing synthetic text generation on three domains—medical, legal, and commercial—we demonstrate an $8\text{-}16\times$ reduction in computation cost across the board to achieve the same privacy and utility. [^2] We allow some baselines an advantage by (incorrectly) reducing their sensitivity by $2$ or $\sqrt{2}$; see §5. Thus, InvisibleInk can produce hundreds of tokens at significantly smaller compute budgets (such as $10\times$ the non-private cost), including in settings where baselines extinguish their privacy budgets within a handful of tokens.
@@
We open source a pip-installable Python package (invink), available at https://github.com/cerai-iitm/invisibleink, to generate differentially private synthetic text using InvisibleInk. See §F for an example of the invink package in action.
@@ 2 Related Work
Generating synthetic data with DP has an extensive literature, particularly over categorical features [21, 22, 23, 24, 25]. Many such approaches rely on the exponential mechanism [26, 27, 28]. Unfortunately, it is cryptographically hard to generate DP synthetic data that preserves all 2-way marginals with polynomial sample complexity [29]; see also [30, Thm. 6.12] and [31, Ch. 9]. Fortunately, public data can provide strong priors over data distributions to partially circumvent these lower bounds. Public data has long been used in private tabular data generation [32, 33, 34, 35]. More recently, pretrained generative foundation models provide strong priors, enabling the DP generation of complex image/text data [36, 37, 38, 39, 40, 41]. In this work, we focus on practical and scalable DP generation of text using LLMs.
We focus on practical and scalable DP text generation. DP text can be generated from DP fine-tuned models [42, 43, 44, 45, 46] or via private inference. [^3] Private inference can be used on pretrained models prompted with private data (our setting) or non-DP fine-tuned models. Both are conceptually identical as they privatize a set of non-private predictions. We focus on the latter as private fine-tuning can be prohibitively expensive, especially with large models [15]. Private inference is possible with various types of LLM access: via a text generation API [47, 41, 46], or white-box access to next-token logits [47, 48, 13]; we focus on the latter. Other DP text manipulation tasks include paraphrasing [49, 50] and next-token prediction [51, 52, 53, 54, 19]. We adapt [19] to text generation as a baseline in §5.
@@
Prior DP text generation approaches diverge significantly from successful decoding algorithms in the NLP literature. Sampling from truncated next-token distributions [58, 16, 59, 60, 61, 62] is strongly backed by qualitative [16], quantitative [17, 18, 63], and theoretical [64] evidence; see the survey [8] for further details. InvisibleInk generates high-utility text by adapting truncated decoding to the DP setting.
The concurrent work of [57] extends the approach of Amin et al. [13] by clustering similar references into a batch. This orthogonal approach can also be integrated with InvisibleInk for synthetic text generation, though clustering may not be applicable to some tasks, such as RAG. [57] also replaces mean aggregation of (clipped) logits with median aggregation, which has a smaller local sensitivity and a tighter ex-post DP guarantee [65], i.e., the DP guarantee is evaluated after the algorithm’s output is observed. In contrast, InvisibleInk provides a stronger standard (ex-ante) DP guarantee, which can be evaluated before the algorithm is run. Similar to [13], the DP guarantee in [57] is data-dependent and shares the same limitations (discussed at the end of §3).
@@ 3 Preliminaries: Differentially Private Text Generation
Language Models & Decoding. An auto-regressive language model $P$ over a vocabulary $V$ defines a distribution $P(x_{t}|{\bm{x}}_{<t})$ over the next token $x_{t}\in V$ in response to a prefix ${\bm{x}}_{<t}:=(x_{1},\ldots,x_{t-1})\in V^{*}$; here, $V^{*}$ denotes the set of all sequences of elements from $V$. These probabilities are typically obtained as the softmax of the next-token logits $\phi(\cdot\,|\,{\bm{x}}_{<t})$ as $P(x_{t}|{\bm{x}}_{<t})\propto\exp\big(\phi(x_{t}|{\bm{x}}_{<t})\big)$.
@@
[…]
@@
where $V_{k}\subset V$ is the set of $k$ tokens with highest logits $\phi(\,\cdot\,|\,{\bm{x}}_{<t})$. See Fig.˜2 for an example.
@@
The quality of non-DP long-form text generation is primarily determined by the decoding algorithm and its hyperparameter ($\tau$ or kk[16, 17, 63]. Typically, these hyperparameters are tuned to balance two competing objectives: reducing (or zeroing out) the probability of generating contextually inappropriate or low-probability tokens, while allowing the selection of moderately probable but contextually appropriate tokens to maintain diversity and coherence. In practice, this typically translates to a temperature $\tau\approx 1$ or slightly smaller, and $k\in[10,100]$ (cf. Fig.˜2).
@@
Figure 2: Left & Center: Illustration of how two common decoding algorithms—temperature rescaling and top-$k$ sampling—reshape the next-token probabilities for (non-private) LLM-based text generation. Right: Heatmap of MAUVE scores [17, 18] of synthetic text generated for the MIMIC-IV-Notes dataset (without using any sensitive references). The best generations (highest MAUVE scores) are obtained at $\tau\approx 1.1$ and $k\approx 100$; InvisibleInk exhibits similar behavior of decoding hyperparameters for private text generation.
@@
In this work, we generate text with a set of privacy-sensitive references ${\bm{R}}=({\bm{r}}_{1},\ldots,{\bm{r}}_{B})\in(V^{*})^{B}$. This setting arises in synthetic text generation, where we instruct an LLM to generate text similar to given references, or in RAG systems where $\bm{R}$} contains retrieved information over confidential knowledge bases. The model output may leak sensitive information from $\bm{R}$} in such cases.
@@
Differential Privacy (DP). DP provides formal protections against such privacy leakage. At a high level, a (randomized) text generation algorithm $\mathcal{A}$ satisfies DP at the example level, if for all adjacent reference sets ${\bm{R}},{\bm{R}}^{\prime}$ (that differ in a single sensitive example), the probability distributions $\mathcal{A}({\bm{q}},{\bm{R}})$ and $\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})$ over responses ${\bm{x}}\in V^{*}$ are “nearly indistinguishable”. A fully specified example-level DP guarantee requires precise notions of adjacency and indistinguishability. We focus on the replace-by-null adjacency, where an adjacent $\bm{R}$$^{\prime}$ is obtained by replacing one element of $\bm{R}$} by the empty string (or vice versa); we denote it as ${\bm{R}}\simeq{\bm{R}}^{\prime}$. This is functionally similar to adding or removing a reference, but simplifies the theoretical technicalities; see [15] or §B for a comparison of adjacency notions. We use the indistinguishability notion of zero-concentrated DP (zCDP) [67]. Let $D_{\alpha}$ denote the $\alpha$-Rényi divergence. Then, an algorithm $\mathcal{A}$ satisfies $\rho$-zCDP if
$D_{\alpha}\big(\mathcal{A}({\bm{q}},{\bm{R}})\|\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})\big)\leq\rho\,\alpha\quad\text{for all}\,\,\alpha>1\,\,\text{ and all adjacent }{\bm{R}}\simeq{\bm{R}}^{\prime}\,.$
@@
Smaller values of $\rho$ mean that $\mathcal{A}({\bm{q}},{\bm{R}})$ and $\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})$ are closer to each other (in terms of the Rényi divergence) and the impact of the differing example between ${\bm{R}},{\bm{R}}^{\prime}$ on the output is low, indicating higher privacy. The zCDP guarantee can be converted into other notions, e.g., $(\varepsilon,\delta)$-DP [68, 69, 70]. Our techniques also translate directly to other notions of adjacency and indistinguishability; see §C.
@@
DP Text Generation via the Exponential Mechanism. Our goal is to design an algorithm $\mathcal{A}({\bm{q}},{\bm{R}})$ to generate text in response to a given query $\bm{q}$} and references $\bm{R}$} with a desired $\rho$-zCDP guarantee. We assume the whitebox setting where model logits $\phi(\,\cdot\,|\,{\bm{x}})$ can be queried for any ${\bm{x}}\in V^{*}$. An increasing number of state-of-the-art general-purpose and reasoning open-weights models including DeepSeek, LLaMA3, Mistral, GPT-OSS allow whitebox access to model logits. This is in contrast to the more restrictive API-access setting, where one inference call returns an entire text sequence [41].
Figure 3: Left two: Histograms of private logits ${\bm{\phi}}_{i}$ and differences from public logits ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ for a synthetic data sample generated from the MIMIC dataset, with 55th and $95$th percentiles shown by the dotted lines. The spread of values for ${\bm{\phi}}_{i}-{\bm{\phi}}_{\mathsf{pub}}$ is significantly smaller (around $10\times$) than that of ${\bm{\phi}}_{i}$. Thus, $\mathsf{DClip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})(y)={\bm{\phi}}_{i}(y)$ for over $95\%$ of all $y\in V$ with $C\approx 1$, while the naive clipping of Amin et al. [13] requires $C\approx 8$. This translates into an $8\times$ gain in computational efficiency. Right two: A small clip norm of $C=1$ (using Amin et al. [13]’s method introduces significant bias resulting in a near-uniform distribution over the vocabulary. In contrast, DClip (see §4) preserves probabilities with minimal distortion even at $C=1$.
@@
We describe the prior state-of-the-art approach of Amin et al. [13], which generates the next token $x_{t}\in V$ via the canonical exponential mechanism [14] over the next-token logits of the given LLM. Given a query $\bm{q}$} and sensitive references $\bm{R}$}, the first step is to obtain the next-token logits ${\bm{\phi}}_{i}=\phi(\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$ for each $i\in[B]$ with an LLM inference call. Second, we clip and aggregate the logits as $\bar{\bm{\phi}}=(1/B)\sum_{i=1}^{B}\mathsf{clip}_{C}({\bm{\phi}}_{i})$, where $\mathsf{clip}_{C}({\bm{\phi}})$ clips each coordinate of ${\bm{\phi}}_{i}$ to lie in [C,C][-C,C]. [^4] The clipping in [13] re-centers the logits (which are invariant to additive shifts). While this is empirically important, we omit the details in our exposition as the underlying mathematical properties do not change. We can then sample a $\rho$-zCDP token $x_{t}\in V$ by sampling [71] from the distribution
@@
where the sensitivity $\mathsf{sens}(\bar{\bm{\phi}})$ measures the maximum change in $\bar{\bm{\phi}}(y)$, when we move to an adjacent set ${\bm{R}}^{\prime}\simeq{\bm{R}}$ of references for any $y\in V$. The clipping and aggregating steps control the sensitivity as $\mathsf{sens}(\bar{\bm{\phi}})=O(C/B)$, [^5] We have $\mathsf{sens}(\bar{\bm{\phi}})=\tfrac{2C}{B}$ (resp. $\tfrac{C}{B}$) under replace-by-null (resp. zero-out) adjacency; cf. §B. In particular, §B.3 corrects technical errors in the privacy analysis of Amin et al. [13] due to imprecise adjacency notions. so that Eq.˜2 reduces to sampling with temperature $\tau=O(C/B\sqrt{\rho})$.
Drawbacks of Prior Work. If the clip norm $C$ is too small, the resulting distortions in the next-token logits boost the probabilities of the unlikely tail tokens (see Fig.˜3), and if $C$ is too large, the temperature $\tau\propto C/B$ becomes too large. Both scenarios lead to degenerate text. Generating non-degenerate text from this method requires taking $C$ sufficiently large (e.g. [13] recommend $C\approx 10$), and $B$ large enough to ensure a temperature of $\tau=O(C/B\sqrt{\rho})\approx 1$.
That is, generating each private token requires a large number $B$ of LLM next-token inference calls, which can be prohibitive if the minimum admissible $B$ is large. In practice, we can typically control $B$, as this is the number of reference documents we use for one synthetic generation, or the number of retrieved references in RAG systems. However, we find in §5 that [13] fails to produce coherent text at a batch size $B$ smaller than 50 or 100, making it prohibitively expensive at scale.
@@
This leaves grid search as the only option to calibrate the noise hyperparameters, i.e., run the algorithm for different hyperparameters and return the output from the experiment whose data-dependent DP guarantee is the closest to the target privacy level. This comes at a significantly increased computation and privacy costs [20]. This important practical issue is neglected by prior works.